[Hallc_running] [New Logentry] Staus update

sanghwa at jlab.org sanghwa at jlab.org
Sun Feb 18 11:55:02 EST 2018

Logentry Text:

Yesterday, we received upto 50uA CW beam back after masking 3P04 BLM (https://logbooks.jlab.org/entry/3533434) and we had smooth running during the day shift until we started having trouble with 3C12 BLM trips. As of now, we are receiving 35uA (stable) beam.

The BLM trips seems to be related to RF instability. The beam delivery was suffered from RF trips besides our BLM trips yesterday. RF system experts will make a plan for addressing the RF instability issues. As a consequence, RF recovery and investigation on the RF instability issues are scheduled tomorrow morning. Beam will be taken away from us around 9am.

While we were suffering from the current limit, we finished the highest Q^2 Run Plan for the CT group, and moved to the intermediate Q^2 kinematics during swing shift. 

The HMS momentum was originally set to -4.278 based on the Run Plan on the Wiki, but it turned out the HMS momentum was entered wrong on the Wiki when the Q^2 was adjusted for the intermediate Q^2 Run Plan. The HMS magnets were reset for p = -4.478 at the beginning of owl shift. The Run Plan on Wiki has been updated with correct HMS momentum (https://logbooks.jlab.org/entry/3533702, https://logbooks.jlab.org/entry/3533733)
We finished C 6% coincidence and singles data taking and now running LH2 coincidence. 1015 good coincidence events have been collected with 6% C (468 events with HMS p =-4.278).

There will be no Daily Running meeting today.


This is a plain text email for clients that cannot display HTML.  The full logentry can be found online at https://logbooks.jlab.org/entry/3533938
-------------- next part --------------
An HTML attachment was scrubbed...
URL: <https://mailman.jlab.org/pipermail/hallc_running/attachments/20180218/b073edf7/attachment.html>

More information about the Hallc_running mailing list